Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Evaporator_2, Page 34 of 35  Not logged in ELOG logo
ID Date Author Target Material Target Oven W.O# Source Status Subject Text Attachments
  29   Tuesday, May 29, 2012, 15:43 Nikita BernierBeOTarget boat #2 FailedSintering BeO pellets #3

    

       
            
  
  28   Monday, May 28, 2012, 11:15 John Wong   ProblemsChiller is leaking water Monday, May 28, 2012, 11:08
It was discovered this morning that
the Evaporator #2 chiller was leaking water
 Chiller_leaky.jpg 
  27   Wednesday, May 23, 2012, 17:45 John WongBeOTarget boat #2 DoneSintering BeO pellets #3

    

       
            
 past_5_hrs.JPGpast_21_hrs.pngpower_lost_2012-05-27_12_0630hr.pngTGHT_600A.pngTGHT@600A_past_12hr.JPG 
  26   Tuesday, May 22, 2012, 16:45 Nikita BernierBeOTarget boat #2 DoneSintering BeO pellets #3

    

       
            
  
  25   Tuesday, May 22, 2012, 15:26 John WongSiC38481-2HP-SISDoneRe-condition the HP-SIS SiC target

    

       
            
  
  24   Tuesday, May 22, 2012, 12:15 Nikita BernierSiC38481-2HP-SISDoneRe-condition the HP-SIS SiC target

    

       
            
 TGHT-VAC1.JPGTBHT-VAC1.JPG 
  23   Friday, May 18, 2012, 13:17 John WongSiC38481-2HP-SISDoneRe-condition the HP-SIS SiC target

    

       
            
  
  22   Friday, May 18, 2012, 12:19 Nikita BernierSiC38481-2HP-SISDoneRe-condition the HP-SIS SiC target

    

       
            
 autoramp1.pngievap2autoheat1.png 
  21   Wednesday, May 16, 2012, 14:04 John Wong   Problems resolvedTBHT cooling line leaky

    

       
            
  
  20   Tuesday, May 15, 2012, 17:06 John Wong   Problems resolvedTBHT power supply faulty Tuesday, May 15, 2012, 17:03
The tube heater power supply current
readback and setpoint are not matching.  Ray
 tubeheater_control.png 
  19   Tuesday, May 15, 2012, 17:03 John Wong   ProblemsTBHT cooling line leaky Tuesday, May 15, 2012, 17:00
It was found that there were water
droplets from the TBHT cooling line.  The
 TBHT_cooling_line_leaky.pdf 
  18   Tuesday, May 15, 2012, 14:00 John Wong   MaintenanceInstalled a new stainless metal heat shield Tuesday, May 15, 2012, 13:45
It was discovered that one of the
insulators that supported the target module
 Evaporator#2_broken_insulator.pdfEvaporator#2_upgrade_.pdf 
  17   Tuesday, May 15, 2012, 12:26 John WongSiC38481-2HP-SISDoneRe-condition the HP-SIS SiC targetTuesday, May 15, 2012, 12:24
The HP-SIS SiC target was made on
Nov 03, 2010.  The target consists of
  
  16   Thursday, April 05, 2012, 16:40 Nikita BernierBeOTarget boat #2 DoneSintering BeO pellets #3Thursday, April 05, 2012, 16:31
Target boat contains 6 BeO
  
  15   Tuesday, April 03, 2012, 09:00 John WongEmptyTa-target boat #2  Sintering Ta-target boat with TaC coatingTuesday, April 03, 2012, 08:54
Yesterday, a new target boat was
sprayed with TaC coating and let dried.
 TGHT_maxcurrent.jpg 
  14   Monday, April 02, 2012, 16:26 Nikita BernierBeOTarget boat #1 DoneSintering new BeO pellets

    

       
            
  
  13   Friday, March 30, 2012, 23:21 John WongBeOTarget boat #1 DoneSintering new BeO pellets

    

       
            
 archive_24hr.jpgTGHT@240A.jpgTGHT@700A.jpgTGHT@750A.jpgTGHT@0A.jpg 
  12   Thursday, March 29, 2012, 22:44 John WongBeOTarget boat #1 DoneSintering new BeO pellets

    

       
            
 Parameter.jpg 
  11   Thursday, March 29, 2012, 19:30 Nikita BernierBeOTarget boat #1 DoneSintering new BeO pellets Thursday, March 29, 2012, 19:17
 
  
  10   Thursday, March 22, 2012, 17:56 Nikita BernierBeOTarget boat #1 DoneSintering BeO pellets

    

       
            
  
ELOG V2.9.2-2455